D 3. Write the coordinates of the points of intersection of the line y = 2x + 3 and the parabola y = Enter answer...
Answer #1

9 ya 72 y = 22 +3 x2= 2x+3 x² - 22-3 =0 a? - 3+-330 212-3) +1(2-3) -0 (x-3) (at) zo - 2=3 orazy 2=3 then ya axp3 52X3+3 -9 22